Circular No. 6278

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1995bd AND 1995an C. Pollas, Observatoire de la Côte d’Azur, reports his discovery, on

technical pan films obtained with the 0.9-m Schmidt telescope by D. Albanese on Dec. 19.93 and 21.96 UT, of an apparent supernova (mag V ∼ 18.8 and 18.2, respectively) in UGC 3151. SN 1995bd is located at α = 4^{h}45^{m}21^{s}.24, δ = +11^{o}04′02′′.5 (equinox 2000.0), which is 23′′.6 east and 1′′.6 south of the galaxy’s center. Nothing is present at this location on the digitized Sky Survey and the original Palomar Survey prints. A faint condensation is visible slightly to the west of SN 1995bd. A nearby star (V ∼ 14) has position end figures 22^{s}.09, 22′′.3. On both films, the type-II SN 1995an is again bright at V ∼ 18.5.

P. Garnavich, A. Riess, R. Kirshner, Center for Astrophysics, report: “Spectra of SN 1995bd obtained by P. Berlind on Dec. 22 (Mt. Hopkins 1.5m Tillinghast telescope) show strong, broad absorption features at 500 and 425 nm, but no clear absorption in the red. The spectrum is similar to those of the peculiar type-Ia supernovae 1991T and 1995ac before maximum. Narrow host-galaxy emission lines show a redshift of 0.016.”

RX J0439.8–6809 AND RX J0550.0–7151 P. C. Schmidtke and A. P. Cowley, Arizona State University, commu-

nicate: “Using data obtained at Cerro Tololo Interamerican Observatory in November, we report the optical identification of two supersoft x-ray sources in the Large Magellanic Cloud. Within the x-ray error circle (Greiner et al. 1994, A.Ap. 281, L61) for RX J0439.8–6809, we find that Star 8 (α = 4^{h}39^{m}49^{s}.6, δ = −68^{o}09′02′′, equinox 2000.0; mean V ∼ 21.8) varies by over half a magnitude on a timescale of less than a day. The star’s exceptionally blue colors of B −V = −0.36 and U −B = −1.22 strongly suggest that it is the optical counterpart of this supersoft x-ray source. If at the distance of the LMC, an absolute magnitude of ∼ +3.5 is indicated. From a ROSAT HRI image, an improved position has been obtained for RX J0550.0–7151, a variable, supersoft x-ray source (α = 5^{h}49^{m}46^{s}.7, δ = −71^{o}49′38′′; Cowley et al. 1993, Ap.J. 418, L63). The source is nearly coincident with a red star (V = 13.53, B − V = +1.45, U − B = +0.86) at α = 5^{h}49^{m}46^{s}.4, δ = −71^{o}49′35′′. Its colors are consistent with a symbiotic or a composite system containing a cool star plus a white dwarf, but spectra are needed.”

1995 December 23 Daniel W. E. Green